# Wav2Small: Distilling Wav2Vec2 to 72K parameters for Low-Resource Speech emotion recognition

### v2

@@ Abstract
Speech Emotion Recognition (SER) needs high computational resources to overcome the challenge of substantial annotator disagreement. Today SER is shifting towards dimensional annotations of arousal, dominance, and valence (A/D/V). Universal metrics as the L2 distance prove unsuitable for evaluating A/D/V accuracy due to non converging consensus of annotator opinions. However, Concordance Correlation Coefficient (CCC) arose as an alternative metric for A/D/V where a model’s output is evaluated to match a whole dataset’s CCC rather than L2 distances of individual audios. Recent studies have shown that wav2vec2.0 / wavLM architectures outputing a float value for each A/D/V dimension achieve today’s State-of-the-art (Sota) CCC on A/D/V. The Wav2Vec2.0 / WavLm family has a high computational footprint, but training small models using human annotations has been unsuccessful [^1] Authors in [1] trained a VGG14 of 4.9M params on MSP Podcast v1.7 using only human annotations. This resulted in very low CCC especially for valence: Their VGG14 achieved only 0.248 valence CCC vs 0.638 achieved by thier Wav2Vec2 (Dawn).. In this paper we use a large Transformer Sota A/D/V model as Teacher/Annotator to train 5 student models: 4 MobileNets and our proposed Wav2Small, using only the Teacher’s A/D/V predictions instead of human annotations. The Teacher model we construct also sets a new Sota on the MSP Podcast dataset. We choose MobileNetV4 / MobileNet-V3 as students, as MobileNet has been designed for fast execution times. We also propose Wav2Small – an architecture designed for minimal parameter number and RAM consumption. Wav2Small with an .onnx (quantized) of only 120KB is a potential solution for A/D/V on hardware with low resources, having only 72K parameters [^2] Convolution followed by BatchNorm then ReLU is a fused op in .onnx. vs 3.12M parameters for MobileNet-V4-Small. 333© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ 1 INTRODUCTION
A need for SER on low resource hardware, drives us to investigate small architectures for A/D/V. Neural Architecture Search has birthed useful architectures for categorical emotions SER [2]. However, Sota A/D/V is dominated by Transformer models, such as Wav2Vec2.0 [1] or WavLM [3], both having a VGG7 audio input extractor. A VGG architecture is chosen because the skip connection of ResNet has high RAM footprint. Conventional components such as Convolution, BatchNorm, ReLU have hardware inter-compatibility [4]. In this paper we use distillation to train 5 small student architectures for A/D/V. We train $3$ MobileNetV4444We use the Top-1 pretrained models from https://huggingface.co/timm/
@@
[4] (-Large, -Medium, -Small), as well as a MobileNetV3-Small [5]555pretrained on Audioset., as well as our proposed architecture named Wav2Small.
@@ 2.2 Data / Augmentation
Distillation for categorical emotions has been used for cross-lingual transfer learning [8] or as a pretrain / finetune paradigm [9]. Here we opt for simultaneous distillation from MSP Podcast and mixed-speaker-speech / non-speech data. We apply always-on, mixup [10] thus clean MSP Podcast speech is never shown to the models. This enables long trainings without overfitting [11]. Extensive work on the most useful labels for SER [12] has revealed that soft labels as produced by Teacher are on-a-par with human annotations [13]. During training we use only the Teacher output A/D/V as ground truth and ignore the MSP Podcast annotations.
@@ 3 Experiments
Authors in [1] trained a VGG14 of 4.9M parameters on MSP Podcast v1.7 annotations. This resulted in very low CCC especially for valence, known as the most difficult dimension to recognize with non transformer models: Their VGG14 achieved only 0.248 valence CCC vs 0.638 achieved by Dawn on MSP Podcast v1.7 test-1 split. Now Fig. 2 shows that our 5 students reach a valence CCC >=>=> =0.37. From the 5 students, MobileNets were pretrained hence they score higher than (non pretrained) Wav2Small. Still, Wav2Small reaches significant CCC, e.g. arousal CCC= 0.66 in MSP Podcast test-1, and = 0.56 in IEMOCAP [18], used as an out of distribution dataset only for evaluation. MobileNetV3-S shows lower CCCs of MobileNetV4-S and also has longer execution: 11ms compared to 5ms as show in Table 1 whose values are measured from .onnx running on Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz. Fig. 3 shows that Wav2Small follows the Teacher arousal and valence faithfully on a language not seen in training. All 5 students have a LogMel layer (incorporated in .onnx via convolution). An extra conv is added to all MobileNets to upscale the 1-channel output of LogMel to 3-channel RGB, needed to feed the pretrained weights.
@@ 3.2 Vectorisation of Time as Channels
The power of Wav2Small comes from a large time resolution (tokens) by having a shallow architecture with 13 chanels. The vectorisation of neighboring tokens into the channel dimension creates 169 ”channels” for the Bayesian pool. The joint processing of the vectorised tokens by the classifier (self.adv(x) in Listing 1.1), boosts Wav2Small to achieve arousal CCC = 0.66 in MSP Podcast, = 0.56 in IEMOCAP: Fig. 2/4 without pretraining. The valence dimension is the most difficult to predict: MobileNetV4-S obtains 0.42 and Wav2Small 0.37 yet in out-of-domain movie as in Fig. 3 Wav2Small does track the output of the Teacher faithfully.
Thinking of the lower valence achieved by all students compared to the Teacher, [19] has shown that Wav2vec2 family of models as they have ASR pretraining, has learned linguistic cues to solve SER. Our 5 students do not posess linguistic cues thus aiming towards a truly unsupervised A/D/V model that is fair for many languages. A model equiped with additional textual features, as the WavLM that won the 2024 SER classification challenge [20] is a future avenue, although under the risk of introducing language/accent unfairness from focusing at words instead of sound’s tonality.
Fig. 1 shows that A/D/V values do separate also the categorical emotions to some extent. One also notices that MobileNetV4-S and Wav2Small obtain an even larger dispersion along the valence axis. Therefore, an auxiliary loss/teacher [^9] https://huggingface.co/3loi/SER-Odyssey-Baseline-WavLM-Categorical should be added to distillation for predicting also categorical emotions. Perhaps via CCC rather than cross-entropy. Observer variability and annotator agreement have been investigated across various annotation schemes [21] with findings revealing higher consensus when annotating A/D/V dimensions compared to annotating one out of six emotion categories. Thus we espy A/D/V to become the major avenue for SER [22] in the future.
@@ 4 CONCLUSION
We propose a 72K parameter architecture named Wav2Small that reaches significant A/D/V score on both MSP Podcast and IEMOCAP datasets, while using only 9 MB RAM compared to 36 MB RAM used by MobileNetV4-S, while having equal MACs. Mobilenet-V4 and MobileNet-V3 reduce the time-sample resolution to a single: 960-dimensional token / 1s audio, whereas Wav2Small provides 250: 169-dimensional-tokens / 1s audio, as its output (before Bayesian pool). Hence Wav2Small is a potential replacement of the expensive input audio extractor of Transformer architectures, such as wav2vec2.0 and WavLM. We also proposed a Teacher for distillation that obtains a new SOTA on MSP Podcast achieving valence CCC = 0.676.

### v3

@@ Abstract
Speech Emotion Recognition (SER) needs high computational resources to overcome the challenge of substantial annotator disagreement. Today SER is shifting towards dimensional annotations of arousal, dominance, and valence (A/D/V). Universal metrics as the L2 distance prove unsuitable for evaluating A/D/V accuracy due to non converging consensus of annotator opinions. However, Concordance Correlation Coefficient (CCC) arose as an alternative metric for A/D/V where a model’s output is evaluated to match a whole dataset’s CCC rather than L2 distances of individual audios. Recent studies have shown that wav2vec2 / wavLM architectures outputing a float value for each A/D/V dimension achieve today’s State-of-the-art (Sota) CCC on A/D/V. The Wav2Vec2.0 / WavLM family has a high computational footprint, but training small models using human annotations has been unsuccessful. In this paper we use a large Transformer Sota A/D/V model as Teacher/Annotator to train 5 student models: 4 MobileNets and our proposed Wav2Small, using only the Teacher’s A/D/V outputs instead of human annotations. The Teacher model we propose also sets a new Sota on the MSP Podcast dataset of valence CCC=0.676. We choose MobileNetV4 / MobileNet-V3 as students, as MobileNet has been designed for fast execution times. We also propose Wav2Small – an architecture designed for minimal parameters and RAM consumption. Wav2Small with an .onnx (quantised) of only 120KB is a potential solution for A/D/V on hardware with low resources, having only 72K parameters vs 3.12M parameters for MobileNet-V4-Small.  111© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ 1 INTRODUCTION
A need for SER on low resource hardware, drives us to investigate small architectures for A/D/V. Neural Architecture Search has birthed useful architectures for categorical emotions SER [2]. However, Sota A/D/V is dominated by Transformer models, such as Wav2Vec2.0 [1] or WavLM [3], both having a VGG7 input feature extractor. This VGG architecture is chosen because the skip connection of ResNet has high RAM footprint. In this paper we apply distillation to train 5 small student architectures for A/D/V. We train $3$ MobileNetV4222We use the Top-1 pretrained models from https://huggingface.co/timm/
@@
[4] (-Large, -Medium, -Small), as well as a MobileNetV3-Small [5]333pretrained on Audioset., as well as our proposed architecture named Wav2Small.
@@ 2.2 Data / Augmentation
Distillation for categorical emotions has been used for cross-lingual transfer learning [8] or as a pretrain / finetune paradigm [9]. Here we opt for simultaneous distillation on MSP Podcast and unlabeled / non-speech data. We apply always-on, mixup [10] thus clean MSP Podcast speech is never shown to the models. This enables long trainings without overfitting [11]. Extensive work on the most useful labels for SER [12] has revealed that soft labels as produced by Teacher provide better of human annotations[13], hence for training we use only the Teacher output A/D/V as ground truth and ignore the MSP Podcast annotations.
@@ 3 Experiments
Authors in [1] trained a CNN14 of 4.9M parameters on MSP Podcast v1.7 human annotations. This resulted in very low CCC especially for valence, known as the most difficult dimension to recognize with non transformer models: Their CNN14 achieved only 0.248 valence CCC vs 0.638 achieved by Dawn on MSP Podcast v1.7 test-1 split. Now Fig. 2 shows that our 5 students reach a valence CCC >=>=> =0.37.
From the 5 students, MobileNets were pretrained and score slighlty higher than (non pretrained) Wav2Small. Still, Wav2Small reaches high CCC, e.g. arousal CCC= 0.66 in MSP Podcast test-1, and CCC=0.56 in IEMOCAP [18], used as an out of distribution dataset only for evaluation. MobileNetV3-S shows lower CCCs of MobileNetV4-S and also has longer execution: 11ms compared to 5ms as show in Table 1 whose values are measured from .onnx running on Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz. All 5 students have a LogMel layer (incorporated in .onnx via convolution). An extra conv is added to all MobileNets to upscale the 1-channel output of LogMel to 3-channel RGB, needed to feed the pretrained weights.
@@ 3.2 Vectorisation of Time as Channels
The power of Wav2Small comes from a large time resolution (tokens) having a shallow architecture with 13 chanels. The vectorisation of neighboring tokens into the channel dimension creates 169 ”channels” for the Bayesian pool. The joint processing of the vectorised tokens by the Bayesian pool / classifier (self.adv(x) in Listing 1.1), boosts Wav2Small to achieve arousal CCC = 0.66 in MSP Podcast, = 0.56 in IEMOCAP: Fig. 2/4 without pretraining. The valence dimension is the most difficult to recognize: MobileNetV4-S obtains CCC=0.42 and Wav2Small CCC=0.37. Nonetheless, Fig. 3 shows that Wav2Small follows the Teacher arousal and valence faithfully on a language not seen in training.
Thinking of the lower valence achieved by all students compared to the Teacher, [19] has shown that Wav2vec2 family of models as they have ASR pretraining, has learned linguistic cues to solve SER. Our 5 students do not posess linguistic cues thus aiming towards a truly unsupervised A/D/V model that is fair for many languages.
A model equipped with textual features, as the WavLM that won the 2024 SER (categorical) challenge [20] is a future avenue for A/D/V, although under the risk of introducing accent unfairness from focusing at words instead of sound’s tonality.
Fig. 1 shows that A/D/V values do separate also the categorical emotions to some extent. One also notices that MobileNetV4-M and Wav2Small obtain an even larger dispersion along the valence axis. Therefore, an auxiliary loss/teacher [^7] https://huggingface.co/3loi/SER-Odyssey-Baseline-WavLM-Categorical should be added to distillation for recognizing also categorical emotions. Perhaps via CCC rather than cross-entropy.
Observer variability and annotator agreement have been investigated across various annotation schemes [21] with findings revealing higher consensus when annotating A/D/V dimensions compared to annotating one out of six emotion categories. Thus we espy A/D/V to become the major avenue for SER [22] in the future.
@@ 4 CONCLUSION
We propose a 72K parameter architecture named Wav2Small that reaches high A/D/V CCC scores on both MSP Podcast and IEMOCAP datasets, while using only 9 MB RAM compared to 36 MB RAM used by MobileNetV4-S, while having equal MACs. Mobilenet-V4 and MobileNet-V3 reduce the time resolution to a single: 960-dimensional token / 1s audio, whereas Wav2Small provides 250: 169-dimensional tokens / 1s audio, as its output (before Bayesian pool). Hence Wav2Small is a potential replacement of the expensive input audio extractor of Transformer architectures, such as wav2vec2 and WavLM. We also proposed a Teacher model for distillation that obtains a new Sota on MSP Podcast achieving valence CCC = 0.676.